a.A saturated solution has:

a. so much solute that it may spontaneously crystallize out

b.enough solute so that some precipitate remains undissolved

c.the ability to absorb easily much more solute

d.the maximum solute that can dissolve in it

b. In which solution will 0.010 M F- remain soluble?

a. 0.020 M Ba2+ (Ksp of BaF2=1.7x10-6)

b. 4.0x10-7 M Ca2+ (Ksp of CaF2=3.9x10-11)

c. 3.0x10-13 M La3+ (Ksp of LaF3=2x10-19)

d. 0.00030 M Pb2+ (Ksp of PbF2=3.6x10-8)

c. Which can be an Arrhenius, Brønsted-Lowry, and Lewis acid?

a. NH3(aq)

b. CO2(g)

c. CH3NH2(aq)

d. H3O+(aq)

d. The Lewis acid-base concept explains better than other concepts the acid-base nature of

a. proton acceptors

b. carboxylic acids

c. alkali metals

d. transition metal ions
Answer #1

a)

Correct option: d. the maximum solute that can dissolve in it

A saturated solution has the maximum amount of solute that can dissolve in it. Any additional solute will remain undissolved.

---

b)

Correct option: d. 0.00030 M Pb2+ (Ksp of PbF2=3.6x10-8)

Ionic product, Qsp of BaF2 = [Ba2+][F-]2 = (0.020)(0.010)2 = 2.0 x 10-6

Ionic product, Qsp of CaF2 = [Ca2+][F-]2 = (4.0x10-7)(0.010)2 = 4.0 x 10-11

Ionic product, Qsp of LaF3 = [La3+][F-]3 = (3.0x10-11)(0.010)3 = 3.0 x 10-17

Ionic product, Qsp of PbF2 = [Pb2+][F-]2 = (0.00030)(0.010)2 = 3.0 x 10-8

For 0.00030 M Pb2+solution Qsp < Ksp. Hence, in 0.00030 M Pb2+, 0.010 M F- will remain soluble